OpenStudy (reverdeparadis):

After 13 minutes of kickboxing class, Jessica has a heart rate of 102. After 15 minutes, she has a heart rate of 106. Assuming that Jessica's heart rate H is a linear function of the time T spent in kickboxing, write H as a linear function of T.

OpenStudy (anonymous):

linear function H(T) will have the form: H = mT + b we want m and b. [taken from a line has the general form y = mx + b] where m is the slope and b is the y intercept. the slope: \[m=\frac{ ΔH }{ ΔT }=\frac{ H_2 - H_1 }{ T_2 - T_1 }= \frac{ 106-102 }{ 15-13 } = \frac{ 4 }{ 2 } = \] to find the y intercept plug in any point (H, T) and the slope you just found. use the first point: (H, T) = (102, 13) and plug it in H = mT + b 102 = 2(13) + b solve for b: b = 76 so the function H(T) is: H = 2T + 76 feel free to ask questions ^_^

OpenStudy (anonymous):

eating supper but ill be back

OpenStudy (skullpatrol):

@reverdeparadis do you HONESTLY think you could of done all of that yourself?

OpenStudy (reverdeparadis):

@skullpatrol Um, excuse me?

OpenStudy (reverdeparadis):

@Euler271 thanks so much! i got it :)

OpenStudy (anonymous):

glad i could help ^_^

OpenStudy (skullpatrol):

@Euler271 you did more than just help, you did the ENTIRE question.